# Local Stability and Region of Attraction Analysis for Neural Network Feedback Systems under Positivity Constraints

### v1

@@ V-A1 Computation of Pre-Activation Logits
Consider a fully connected FFNN described in 16. The pre-activation bounds are computed recursively using interval arithmetic as follows.
Assuming the input $y\in\Gamma\subseteq\mathbb{R}^{n_{0}}$ is bounded componentwise by $\underline{y}\leq y\leq\overline{y}$, the input bounds propagate through the first layer using:
@@
$\underline{\nu}_{j}^{(1)}=\sum_{k=1}^{n_{0}}\min\left\{W^{(1)}_{jk}\underline{y}_{k},W^{(1)}_{jk}\overline{y}_{k}\right\},$ (19)
$\overline{\nu}_{j}^{(1)}=\sum_{k=1}^{n_{0}}\max\left\{W^{(1)}_{jk}\underline{y}_{k},W^{(1)}_{jk}\overline{y}_{k}\right\}.$ (20)
@@
$\underline{z}^{(i-1)}$ $=\phi^{(i-1)}(\underline{\nu}^{(i-1)}),\quad\overline{z}^{(i-1)}=\phi^{(i-1)}(\overline{\nu}^{(i-1)}),$ (21)
@@
These recursive bounds can be computed efficiently, and they enable us to define the domain over which linear relaxations of activation functions are constructed. The resulting pre-activation intervals $[\underline{\nu}^{(i)},\overline{\nu}^{(i)}]$ are then used to form upper and lower linear envelopes for each activation function.
(a) Positive $\nu$: $\beta=\frac{\tanh(\overline{\nu})}{\overline{\nu}}$, $\alpha=\frac{\tanh(\underline{\nu})}{\underline{\nu}}$, giving $\beta\,\underline{\nu}\leq\tanh(\nu)\leq\alpha\,\overline{\nu}$.
(b) Negative $\nu$: $\beta=\frac{\tanh(\underline{\nu})}{\underline{\nu}}$, $\alpha=\frac{\tanh(\overline{\nu})}{\overline{\nu}}$, giving $\alpha\,\underline{\nu}\leq\tanh(\nu)\leq\beta\,\overline{\nu}$.
(c) Unstable $\nu$: $\alpha_{1}=\alpha_{2}=\frac{d}{d\nu}\tanh(\nu)=1$, $\beta_{1}=\frac{\tanh(\underline{\nu})}{\underline{\nu}}$, $\beta_{2}=\frac{\tanh(\overline{\nu})}{\overline{\nu}}$, so that $\alpha_{1}\,\underline{\nu}\leq\tanh(\nu)\leq\alpha_{2}\,\overline{\nu}$.
Figure 2: Illustration of the linear relaxation of $\phi(\nu)=\tanh(\nu)$ over different pre-activation intervals.
@@
V-A2 linear relaxation of activation functions
@@
$\begin{cases}\beta_{j}^{(i)}\underline{\nu}_{j}^{(i)}\leq\phi(\nu_{j}^{(i)})\leq\alpha_{j}^{(i)}\overline{\nu}_{j}^{(i)}&\text{if }\underline{\nu}_{j}^{(i)}\geq 0,\ \overline{\nu}_{j}^{(i)}>0,\\ \alpha_{j}^{(i)}\underline{\nu}_{j}^{(i)}\leq\phi(\nu_{j}^{(i)})\leq\beta_{j}^{(i)}\overline{\nu}_{j}^{(i)}&\text{if }\underline{\nu}_{j}^{(i)}<0,\ \overline{\nu}_{j}^{(i)}\leq 0,\\ -|\alpha_{1,j}^{(i)}\underline{\nu}_{j}^{(i)}|\leq\phi(\nu_{j}^{(i)})\leq|\alpha_{2,j}^{(i)}\overline{\nu}_{j}^{(i)}|&\text{if }\underline{\nu}_{j}^{(i)}<0,\ \overline{\nu}_{j}^{(i)}>0.\end{cases}$ (26)
@@
$\underline{\nu}_{j}^{(i)}=L^{(i)}y\quad\overline{\nu}_{j}^{(i)}=U^{(i)}y,$ (27)
respectively. For the third case (crossing zero), we apply the Cauchy–Schwarz inequality:
$-|\alpha_{1,j}^{(i)}|\,|\underline{\nu}_{j}^{(i)}|\leq\phi(\nu_{j}^{(i)})\leq|\alpha_{2,j}^{(i)}|\,|\overline{\nu}_{j}^{(i)}|.$ (28)
Now, we replace $\underline{\nu}_{j}^{(i)}$ and ${\overline{\nu}}_{j}^{(i)}$ by $L^{(i)}\,y$ and $U^{(i)}\,y$ respectively, and since $y\geq 0$ for positive systems, we can move the vector $y$ outside the absolute value, yielding:
$-|\alpha_{1,j}^{(i)}|\,|L^{(i)}|y\leq\phi(\nu_{j}^{(i)})\leq|\alpha_{2,j}^{(i)}|\,|U^{(i)}|y.$ (29)
We now combine equations 26, 27, and (29) to form a unified equation. Let $D^{(i)}_{\text{low}},D^{(i)}_{\text{up}}\in\mathbb{R}^{n_{i}\times n_{i}}$ be diagonal matrices containing the slope parameters and define adjusted slope matrices $\hat{L}^{(i)},\hat{U}^{(i)}\in\mathbb{R}^{n_{i}\times n_{0}}$. The post-activation output is sector bounded as:
$D^{(i)}_{\text{low}}\,\hat{L}^{(i)}y\leq\phi(\nu^{(i)})\leq D^{(i)}_{\text{up}}\,\hat{U}^{(i)}y,$ (30)
@@
$\begin{cases}D^{(i)}_{\text{low},jj}=\beta_{j}^{(i)},\quad D^{(i)}_{\text{up},jj}=\alpha_{j}^{(i)},\\ \hat{L}_{j,:}^{(i)}=L_{j,:}^{(i)},\quad\hat{U}_{j,:}^{(i)}=U_{j,:}^{(i)}&\text{if }\underline{\nu}_{j}^{(i)}\geq 0,\ \overline{\nu}_{j}^{(i)}>0,\\[4.30554pt] D^{(i)}_{\text{low},jj}=\alpha_{j}^{(i)},\quad D^{(i)}_{\text{up},jj}=\beta_{j}^{(i)},\\ \hat{L}_{j,:}^{(i)}=L_{j,:}^{(i)},\quad\hat{U}_{j,:}^{(i)}=U_{j,:}^{(i)}&\text{if }\underline{\nu}_{j}^{(i)}<0,\ \overline{\nu}_{j}^{(i)}\leq 0,\\[4.30554pt] D^{(i)}_{\text{low},jj}=-|\alpha_{1,j}^{(i)}|,\quad D^{(i)}_{\text{up},jj}=|\alpha_{2,j}^{(i)}|,\\ \hat{L}^{(i)}_{j,:}=|L^{(i)}_{j,:}|,\quad\hat{U}^{(i)}_{j,:}=|U^{(i)}_{j,:}|&\text{if }\underline{\nu}_{j}^{(i)}<0,\ \overline{\nu}_{j}^{(i)}>0,\end{cases}$ (31)
where $L_{j,:}$ denotes the $j$-th row of matrix $L.$
@@ V-B Propagation of Sector Bounds through Weight Matrices
$l^{(i)}y\leq\omega^{(i)}\leq u^{(i)}y,$ (32)
with
$l^{(i)}=D^{(i)}_{\text{low}}\,\hat{L}^{(i)},\quad u^{(i)}=D^{(i)}_{\text{up}}\,\hat{U}^{(i)}.$
@@ V-C Sector Bound for the Entire Neural Network
Consider a fully connected FFNN as in (16) with sector bounded monotonic activation functions. Given the set $\Gamma$, then the sector bound for the network mapping $\mathrm{NN}(y)$ can be calculated as:
$\gamma_{1}\,y\leq\mathrm{NN}(y)\leq\gamma_{2}\,y,\quad\forall y\in\Gamma,$ (34)
$\gamma_{1}=L^{(q+1)}\left(\prod_{i=2}^{q}\underline{D}^{(i)}\hat{L}^{(i)}\right)W^{(1)},$
$\gamma_{2}=U^{(q+1)}\left(\prod_{i=2}^{q}\overline{D}^{(i)}\hat{U}^{(i)}\right)W^{(1)},$
and the matrices $\underline{D}^{(i)},\overline{D}^{(i)},\hat{L}^{(i)},\hat{U}^{(i)}$ are defined according to the activation relaxation scheme described in (31).
@@
This result follows directly from Theorem 2 and Theorem 4. Furthermore, in the scalar case (i.e., when the Lur’e system is MISO), Lemma 3 can be applied to estimate the region of attraction. To do so, one computes the maximal element of $\Gamma$, and then applies the analytical formula in (5) to determine the set of stable initial conditions.
@@ VI Numerical Example
$A=\begin{bmatrix}-7&5\\ 6&1\end{bmatrix},\quad B=\begin{bmatrix}1\\ 2\end{bmatrix},\quad C=\begin{bmatrix}1&1\end{bmatrix}.$
[…]
@@
We now apply the Lyapunov-based approach by solving the matrix inequality (10). This yields a quadratic Lyapunov function $V(x)=x^{\top}P_{\text{lm}}x$ with $P_{\text{lm}}=\begin{bmatrix}0.1235&0.1323\\ 0.1323&0.3919\end{bmatrix}.$ Using Algorithm 1, we estimate the largest sublevel set satisfying $\dot{V}(x)<0$, resulting in the ROA $x^{\top}P_{\text{lm}}x\leq 32.$ Figure 4a visualizes this approach by showing sublevel sets that remain in the negative $\dot{V}$ area, and Figure 4b shows random trajectories starting inside the estimated ROA.
@@
Figure 5: Local sector bounds calculated for given $y\in\Gamma$ using equation (34). The title of each subplot defines the $\Gamma$ set and the corresponding $[\gamma_{1},\gamma_{2}]$.
@@
As a benchmark, we compare our method against one of the best approaches in the literature [2], which uses IQCs to bound NN nonlinearities. This method produces a Lyapunov function $V(x)=x^{\top}P_{\text{qc}}x$, where $P_{\text{qc}}=\begin{bmatrix}0.1675&-0.0224\\ -0.0224&0.0668\end{bmatrix},$ and yields a conservative ROA of $x^{\top}P_{\text{qc}}x\leq 1.$

### v2

@@ V-A1 Computation of pre-Activation logits
Consider a fully connected FFNN described in (16). The pre-activation bounds are computed recursively using interval arithmetic as follows.
Assuming the input $y\in\Gamma\subseteq\mathbb{R}^{n_{0}}$ is bounded component-wise by $\underline{y}\leq y\leq\overline{y}$, the input bounds propagate through the first layer using:
@@
$\small\underline{\nu}_{j}^{(1)}=\sum_{k=1}^{n_{0}}\min\left\{W^{(1)}_{jk}\underline{y}_{k},W^{(1)}_{jk}\overline{y}_{k}\right\},$ (19)
$\small\overline{\nu}_{j}^{(1)}=\sum_{k=1}^{n_{0}}\max\left\{W^{(1)}_{jk}\underline{y}_{k},W^{(1)}_{jk}\overline{y}_{k}\right\}.$ (20)
@@
$\small\underline{z}^{(i-1)}$ $=\phi^{(i-1)}(\underline{\nu}^{(i-1)}),\quad\overline{z}^{(i-1)}=\phi^{(i-1)}(\overline{\nu}^{(i-1)}),$ (21)
@@
These recursive bounds can be computed efficiently, and they enable us to define the domain over which linear relaxations of activation functions are constructed. The resulting pre-activation intervals $\underline{\nu}^{(i)},\overline{\nu}^{(i)}]$ are then used to form upper and lower linear envelopes for each activation function.
(a) Positive $\nu$: $\beta=\frac{\tanh(\overline{\nu})}{\overline{\nu}},\alpha=\frac{\tanh(\underline{\nu})}{\underline{\nu}},$
$\beta\underline{\nu}\leq\tanh(\nu)\leq\alpha\overline{\nu}$
(b) Negative $\nu$: $\beta=\frac{\tanh(\underline{\nu})}{\underline{\nu}},\alpha=\frac{\tanh(\overline{\nu})}{\overline{\nu}},$
$\alpha\underline{\nu}\leq\tanh(\nu)\leq\beta\overline{\nu}$
(c) Unstable $\nu$: $\alpha_{1}=\alpha_{2}=\frac{d\tanh(\nu)}{d(\nu)}=1,$
$\beta_{1}=\frac{\tanh(\underline{\nu})}{\underline{\nu}},\beta_{2}=\frac{\tanh(\overline{\nu})}{\overline{\nu}},$
$\alpha_{1}\underline{\nu}\leq\tanh{(\nu)}\leq\alpha_{2}\overline{\nu}$
Figure 2: Illustration of linear relaxation of $\phi=\tanh(\nu)$ under different intervals for $\nu$.
@@
V-A2 Sector relaxation of activation functions
@@
$\small\begin{cases}\beta_{j}^{(i)}\underline{\nu}_{j}^{(i)}\leq\phi(\nu_{j}^{(i)})\leq\alpha_{j}^{(i)}\overline{\nu}_{j}^{(i)}&\text{if }\underline{\nu}_{j}^{(i)}\geq 0,\ \overline{\nu}_{j}^{(i)}>0,\\ \alpha_{j}^{(i)}\underline{\nu}_{j}^{(i)}\leq\phi(\nu_{j}^{(i)})\leq\beta_{j}^{(i)}\overline{\nu}_{j}^{(i)}&\text{if }\underline{\nu}_{j}^{(i)}<0,\ \overline{\nu}_{j}^{(i)}\leq 0,\\ -|\alpha_{1,j}^{(i)}\underline{\nu}_{j}^{(i)}|\leq\phi(\nu_{j}^{(i)})\leq|\alpha_{2,j}^{(i)}\overline{\nu}_{j}^{(i)}|&\text{if }\underline{\nu}_{j}^{(i)}<0,\ \overline{\nu}_{j}^{(i)}>0.\end{cases}$ (26)
@@
$\small\underline{\nu}_{j}^{(i)}=L_{j,:}^{(i)}y\quad\overline{\nu}_{j}^{(i)}=U_{j,:}^{(i)}y,$ (27)
respectively, where $L_{j,:}$ denotes the $j$-th row of matrix $L$. For the third case (crossing zero), we apply the Cauchy–Schwarz inequality:
$\small-|\alpha_{1,j}^{(i)}|\,|\underline{\nu}_{j}^{(i)}|\leq\phi(\nu_{j}^{(i)})\leq|\alpha_{2,j}^{(i)}|\,|\overline{\nu}_{j}^{(i)}|.$ (28)
Now, we replace $\underline{\nu}_{j}^{(i)}$ and ${\overline{\nu}}_{j}^{(i)}$ by $L_{j,:}^{(i)}\,y$ and $U_{j,:}^{(i)}\,y$ respectively, and since $y\geq 0$ for positive systems, we can move the vector $y$ outside the absolute value, yielding:
$\small-|\alpha_{1,j}^{(i)}|\,|L_{j,:}^{(i)}|y\leq\phi(\nu_{j}^{(i)})\leq|\alpha_{2,j}^{(i)}|\,|U_{j,:}^{(i)}|y.$ (29)
We now combine equations (26), 27, and (29) to form a unified equation. Let $D^{(i)}_{\text{low}},D^{(i)}_{\text{up}}\in\mathbb{R}^{n_{i}\times n_{i}}$ be diagonal matrices containing the slope parameters and define adjusted slope matrices $\hat{L}^{(i)},\hat{U}^{(i)}\in\mathbb{R}^{n_{i}\times n_{0}}$. The post-activation output is sector bounded as:
$\small D^{(i)}_{\text{low}}\,\hat{L}^{(i)}y\leq\phi(\nu^{(i)})\leq D^{(i)}_{\text{up}}\,\hat{U}^{(i)}y,$ (30)
@@
$\small\begin{cases}D^{(i)}_{\text{low},jj}=\beta_{j}^{(i)},\qquad D^{(i)}_{\text{up},jj}=\alpha_{j}^{(i)},\\ \hat{L}_{j,:}^{(i)}=L_{j,:}^{(i)},\qquad\hat{U}_{j,:}^{(i)}=U_{j,:}^{(i)}&\text{if }\underline{\nu}_{j}^{(i)}\geq 0,\ \overline{\nu}_{j}^{(i)}>0,\\[3.87498pt] D^{(i)}_{\text{low},jj}=\alpha_{j}^{(i)},\qquad D^{(i)}_{\text{up},jj}=\beta_{j}^{(i)},\\ \hat{L}_{j,:}^{(i)}=L_{j,:}^{(i)},\qquad\hat{U}_{j,:}^{(i)}=U_{j,:}^{(i)}&\text{if }\underline{\nu}_{j}^{(i)}<0,\ \overline{\nu}_{j}^{(i)}\leq 0,\\[3.87498pt] D^{(i)}_{\text{low},jj}=-|\alpha_{1,j}^{(i)}|,\qquad D^{(i)}_{\text{up},jj}=|\alpha_{2,j}^{(i)}|,\\ \hat{L}^{(i)}_{j,:}=|L^{(i)}_{j,:}|,\qquad\hat{U}^{(i)}_{j,:}=|U^{(i)}_{j,:}|&\text{if }\underline{\nu}_{j}^{(i)}<0,\ \overline{\nu}_{j}^{(i)}>0.\end{cases}$ (31)
@@ V-B Propagation of Sector Bounds through Weight Matrices
$l^{(i)}y\leq\omega^{(i)}\leq u^{(i)}y,\quad l^{(i)}=D^{(i)}_{\text{low}}\,\hat{L}^{(i)},\quad u^{(i)}=D^{(i)}_{\text{up}}\,\hat{U}^{(i)}.$
@@ V-C Sector Bound for the Entire Neural Network
Consider an FFNN as in (16) $\pi:\mathbb{R}_{+}^{p}\to\mathbb{R}^{m}$ with sector bounded activation functions. Given the set $\Gamma$, then the sector bound for the network mapping $\mathrm{NN}(y)$ can be calculated as:
$\gamma_{1}\,y\leq\mathrm{NN}(y)\leq\gamma_{2}\,y,\qquad\forall y\in\Gamma,$ (33)
$\gamma_{1}=L^{(q+1)}\left(\prod_{i=2}^{q}D_{\text{low}}^{(i)}\hat{L}^{(i)}\right)W^{(1)},$
$\gamma_{2}=U^{(q+1)}\left(\prod_{i=2}^{q}D_{\text{up}}^{(i)}\hat{U}^{(i)}\right)W^{(1)},$
and the matrices $D_{\text{low}}^{(i)},D_{\text{up}}^{(i)},\hat{L}^{(i)},\hat{U}^{(i)}$ are defined according to the activation relaxation scheme described in (31).
@@
This result follows directly from Theorem 2 and Theorem 4. Furthermore, in the scalar case (i.e., when the Lur’e system is MISO), Lemma 3 can be applied to estimate the region of attraction. To do so, one computes the upper bound of $\Gamma$, and then applies the analytical formula in (5) to determine the set of stable initial conditions.
@@ VI Numerical Example
$\small A=\begin{bmatrix}-7&5\\ 6&1\end{bmatrix},\quad B=\begin{bmatrix}1\\ 2\end{bmatrix},\quad C=\begin{bmatrix}1&1\end{bmatrix}$.
Using parametric analysis, we compute the minimal lower sector $\Sigma_{1}=-3$ such that the matrix $A+B\Sigma_{1}C$ becomes Metzler, and the maximal upper sector $\Sigma_{2}=-1.276$ such that $A+B\Sigma_{2}C$ is Hurwitz. These values define the stable sector interval $\Sigma_{1},\Sigma_{2}]$ for the application of Aizerman-type results. A PID controller with gains P=1.25P=-1.25, I=0.1I=-0.1, and D=0.1D=-0.1 is used as a benchmark for generating training data for a stabilizing NN controller. We collect closed loop input-output trajectory data by randomly initializing the system near $x_{0}=[2,\ 2]^{\top}$. An FFNN is trained to approximate the stabilizing PID controller. Figure 3a shows the resulting NN and its input-output relation relative to the stable sector bounds $\Sigma_{1},\Sigma_{2}]$. It is evident that the trained NN is $\Gamma$-sector bounded within the interval. Numerically, we identify the $\Gamma$ set and its maximal element to be $\overline{y}=12.2$.
@@
We now apply the Lyapunov-based approach by solving the matrix inequality (10). This yields a quadratic Lyapunov function
$\small V(x)=x^{\top}P_{\text{lm}}x,\textrm{ with }P_{\text{lm}}=\begin{bmatrix}0.1235&0.1323\\ 0.1323&0.3919\end{bmatrix}$
Using Algorithm 1, we estimate the largest sublevel set satisfying $\dot{V}(x)<0$, resulting in the ROA $x^{\top}P_{\text{lm}}x\leq 32.$ Figure 4a visualizes this approach by showing sublevel sets that remain in the negative $\dot{V}$ area, and Figure 4b shows random trajectories starting inside the estimated ROA.
@@
Figure 5: Local sector bounds calculated for given $y\in\Gamma$ using equation (33). The title of each subplot defines the $\Gamma$ set and the corresponding $\gamma_{1},\gamma_{2}]$.
@@
As a benchmark, we compare our method against one of the best approaches in the literature [4], which uses IQCs to bound NN nonlinearities. This method produces a Lyapunov function
$\small V(x)=x^{\top}P_{\text{qc}}x,\textrm{ with }P_{\text{qc}}=\begin{bmatrix}0.1675&-0.0224\\ -0.0224&0.0668\end{bmatrix}$,
and yields a conservative ROA of $x^{\top}P_{\text{qc}}x\leq 1.$